XYZ Co. has the following financial information:

Debt: 25,000 bonds outstanding with a face value of $1,000. The bonds currently trade at 91% of par and have 10 years to maturity. The coupon rate equals 3%, and the bonds make semi-annual interest payments.

Preferred stock: 300,000 shares of preferred stock outstanding; currently trading for $153 per share and it pays a dividend of $6.40 per share every year.

Common stock: 1,000,000 shares of common stock outstanding; currently trading for $85 per share. Beta equals 0.88.

Market and firm information: The expected return on the market is 12%, the risk free rate is 3%, tax rate is 21%

a) What is the weight of common stock in the capital structure? Enter percentages as decimals and round to 4 decimals

b) What is the weight of debt in the capital structure? Enter percentages as decimals and round to 4 decimals

c) What is the before tax cost of debt? Enter percentages as decimals and round to 4 decimals

d) What is the after tax cost of debt? Enter percentages as decimals and round to 4 decimals

e) What is the cost of preferred stock? Enter percentages as decimals and round to 4 decimals

f) What is the cost of common stock? Enter percentages as decimals and round to 4 decimals

g) What is the weighted average cost of capital? Enter percentages as decimals and round to 4 decimals
Answer #1

a. Value of Common Stock =Number of Shares*Share price =1,000,000*85 =85000000
Value of Debt =Par Value*91%*Number of Bonds =1000*91%*25000 =22750000
Value of Preferred Stock =Preferred Share Price*No of shares =153*300000=45900000
Total Value =85000000+22750000+45900000 =153,650,000
Weight of common stock =85000000/153,650,000 =0.5532

b. Weight of Debt =22,750,000/153,650,000 =0.1481

c.Semi annual coupon =3%*1000/2 =15
Number of Periods =2*10 =20
Price =910
Par Value =1000
YTM or cost of debt using excel formula =2*RATE(20,15,-910,1000) = 0.0411

Before tax cost of debt =0.0411

d. After tax cost of debt =0.0411*(1-21%) =0.0324